OpenStudy (anonymous):

prove 1+sinx=2cos^2x

OpenStudy (helder_edwin):

it is not true! if x=0 then the left side is \[\large 1+\sin0=1+0=1 \] and the right side \[\large 2\cos^20=2(1)^2=2 \]

OpenStudy (anonymous):

this is great! how do we know for sure?

OpenStudy (helder_edwin):

what i wrote is counterexample. it definitely shows it is not true.

OpenStudy (helder_edwin):

*is a counterexample.

OpenStudy (anonymous):

yes, you are right. would we be able to prove it by taking the arc of it?

OpenStudy (helder_edwin):

to prove it is not true?

OpenStudy (anonymous):

no, to prove it is true

OpenStudy (helder_edwin):

but it is not true. if it were true, using inverse trig functions usually is not easy 'cause they don't have any useful properties.

OpenStudy (anonymous):

okay! i understand now. this is very helpful thank you so much. im definitely becoming a fan. are you a math teacher? i also have other questions

OpenStudy (helder_edwin):

glad to help

OpenStudy (anonymous):

solve for x : tan^2x-tanx-2=0

OpenStudy (helder_edwin):

u have \[\large \tan^2x-\tan x-2=0 \] use the substitution \(u=\tan x\) so the equation becomes a common quadratic equation.

OpenStudy (anonymous):

okay, so what would that give us?

OpenStudy (helder_edwin):

\[\large u^2-u-2=0 \]

OpenStudy (anonymous):

yes but that does not solve anything where would be go from here

OpenStudy (helder_edwin):

did u solve it?

OpenStudy (helder_edwin):

what did u get?

OpenStudy (anonymous):

u^2-u=2

OpenStudy (helder_edwin):

i mean find the roots, using the quadratic formula or factoring it.

OpenStudy (anonymous):

u(u-1)

OpenStudy (helder_edwin):

do u know the quadratic formula?

OpenStudy (anonymous):

yes -b + or - the sqaure root of b^2-4ac over 2a but i already factored it so what is the need for the formula?

OpenStudy (helder_edwin):

u didn't factor it correctly

OpenStudy (helder_edwin):

\[\large 0=u^2-u-2=(u-2)(u+1)\]

OpenStudy (anonymous):

so you would take each one and set it equal to 0 giving u=2 and u=-1

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

or tanx= 2 and tan x=-1

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

tanx of 2 is extraneous but tanx=-1 can be written as 3pi/4 and 7pi/4

OpenStudy (anonymous):

that is what i had as my original answer i wasnt sure if it is right though

OpenStudy (helder_edwin):

yes, the solutions r right

OpenStudy (helder_edwin):

but \[\large 2=u=\tan x \] would give you x=63.43ยบ (using a calculator)

OpenStudy (anonymous):

yes but the answer should translate to radian measure

OpenStudy (helder_edwin):

in radians it would be x=1.107 (using a calculator)

OpenStudy (anonymous):

so it would be close to pi/3

OpenStudy (helder_edwin):

yes. close to pi/3. but if u have to use tables (not allowed to use calculators) or elegant answers such as \(3\pi/4\), then u=2 can be "dismissed". You should ask your teacher.

OpenStudy (anonymous):

yes making it extraneous. perfect. you are great

OpenStudy (anonymous):

i hope i am not bothering you this homework is very hard. i have a couple of more questions

OpenStudy (helder_edwin):

i am fine. go ahead.

OpenStudy (anonymous):

solve for x: 2sinxcosx=sin2x

OpenStudy (anonymous):

would we rearrange the equation to 2sinxcosx-sinx2x=0

OpenStudy (helder_edwin):

it is \[\large 2\sin x\cos x-\sin x\cdot 2x=0 \] ?

OpenStudy (anonymous):

no the original problem is \[2sinxcosx=\sin2x\]

OpenStudy (helder_edwin):

this is an identity (always true)

OpenStudy (anonymous):

yes but it is asking to solve for x not prove

OpenStudy (anonymous):

the identity is 2sinAcosA=Sin2A

OpenStudy (helder_edwin):

yeah, but u don't solve an identity. any x will solve it.

OpenStudy (anonymous):

yes you are right!

OpenStudy (anonymous):

verify: \[(2+\tan ^{2}x \div \sec ^{2}x) -1\]

OpenStudy (helder_edwin):

aren't u missing an equal sign?

OpenStudy (anonymous):

no it says to verify

OpenStudy (helder_edwin):

but it is not a valid proposition. there's nothing to verify, at least from my point of view. maybe there's a typo in the question.

OpenStudy (anonymous):

no its like you have to use trigonometric identities to prove it, or, to simply. lets say it is to simplify

OpenStudy (helder_edwin):

ok \[\large \frac{2+\tan^2x}{\sec^2x}-1=\frac{2+\tan^2x-\sec^2x}{\sec^2x} \]

OpenStudy (helder_edwin):

now since \[\large \tan^2x+1=\sec^2x \]

OpenStudy (helder_edwin):

then \[\large =\frac{2-1}{\sec^2x}=\frac{1}{\sec^2x}=\cos^2x\]

OpenStudy (helder_edwin):

and that's it, I think

OpenStudy (anonymous):

ok! i have a few more similiar to this one simplify \[\tan ^{2}x (\sec ^{2}x \div \csc ^{2}x)\]

OpenStudy (helder_edwin):

use the definitions of sec and csc \[\large \sec x=1/\cos x\qquad\csc x=1/\sin x \]

OpenStudy (helder_edwin):

sorry, but here it is late. I gotta get some rest.

OpenStudy (helder_edwin):

see u some other time.

OpenStudy (anonymous):

thank you very much for your help

OpenStudy (helder_edwin):

u r welcome. bye